# daVinci-Env: Open SWE Environment Synthesis at Scale

### 13 Mar 2026 v1

@@ 1 Introduction
However, constructing high-quality and diverse executable environments at scale remains a critical bottleneck. While recent open-source efforts such as SWE-rebench (Badertdinov et al., 2025), SWE-Universe (Chen et al., 2026), and SWE-Factory (Guo et al., 2026) have made progress toward automation, the resource barrier is prohibitive: the computational and infrastructure costs of generating validated environments at scale remain extraordinarily high, effectively excluding most academic research groups and creating a stark divide between industrial solutions, which achieve scale but remain opaque with unreleased infrastructure (Chen et al., 2026; Liu et al., 2025a), and open-source alternatives that remain limited in both scale and repository diversity.
@@ 2.1 Environment Synthesis
The construction of executable environments for agents has become a central infrastructure challenge. SWE-bench (Jimenez et al., 2023) pioneered this direction by curating a benchmark of real GitHub issues paired with pull requests, where each task instance is embedded in a Docker-based repository snapshot with executable test suites that serve as evaluation oracles. To overcome this bottleneck, several concurrent efforts have emerged to automate large-scale environment generation. SWE-rebench (Badertdinov et al., 2025) introduces a scalable pipeline that replicates the SWE-bench construction process across a broader set of repositories, aiming to generate thousands of additional task instances with executable test environments. SWE-Universe (Chen et al., 2026) takes a complementary approach by systematically crawling and filtering GitHub repositories to produce a diverse universe of candidate environments. SWE-Factory (Guo et al., 2026) further automates the end-to-end pipeline from repository selection to Dockerfile synthesis and test harness generation. SWE-World (Sun et al., 2026)world proposes an orthogonal direction by replacing physical Docker execution with learned surrogate models trained on agent-environment interaction data, eliminating the resource-intensive costs of Docker environment maintenance while preserving the agent-environment feedback loop.
@@ 3.4 Dockerfile Construction
The Dockerfile agent is responsible for generating a containerized environment for each task. During the pilot study, we identified two recurring failure modes: (1) network instability during environment construction, where generic base images require downloading Python and dependencies at build time, leading to frequent timeouts; and (2) redundant rebuilds, where unchanged base layers are reconstructed from scratch on every iteration. These inefficiencies become particularly costly at scale; therefore, we equip the Dockerfile agent with the following strategies.
@@
The Dockerfile agent will receive repository exploration agent’s findings (e.g., special dependencies from README.md) as additional input, allowing the agent to make more informed initial decisions, and it will operate iteratively to construct the Dockerfile. If the final test execution fails, the Dockerfile agent will also receive the feedback from the test analysis agent and refine its output in subsequent attempts.
@@ 4.2 Main Results
Table 3 presents the comparison of OpenSWE with representative SFT-based methods on SWE-Bench Verified.
@@
OpenSWE-32B achieves a resolution rate of 62.4%, surpassing all existing 32B-scale SFT methods. Compared to the strongest Qwen2.5-Coder-32B baseline SWE-Master-32B and SWE-Master-32B-RL. OpenSWE-32B improves by 4.6% while using a non-Coder base model, demonstrating that high-quality environment data can compensate for domain-specific pretraining. At the 72B scale, OpenSWE-72B reaches 66.0%, outperforming daVinci-Dev-72B by 7.5%. Both the 32B result and 72B result prove the effectiveness of OpenSWE.
@@ 4.4 Impact of Environment Source
To understand how the choice of environment synthesis pipeline affects downstream agent performance, we train identical models on environments from different sources and evaluate under the same protocol. Table 4 reports the results.

### 16 Mar 2026 v2

@@ 1 Introduction
However, constructing high-quality and diverse executable environments at scale remains a critical bottleneck. While recent open-source efforts such as SWE-rebench (Badertdinov et al., 2025), SWE-Universe (Chen et al., 2026b), and SWE-Factory (Guo et al., 2026) have made progress toward automation, the resource barrier is prohibitive: the computational and infrastructure costs of generating validated environments at scale remain extraordinarily high, effectively excluding most academic research groups and creating a stark divide between industrial solutions, which achieve scale but remain opaque with unreleased infrastructure (Chen et al., 2026b; Liu et al., 2025a), and open-source alternatives that remain limited in both scale and repository diversity.
@@ 2.1 Environment Synthesis
The construction of executable environments for agents has become a central infrastructure challenge. SWE-bench (Jimenez et al., 2023) pioneered this direction by curating a benchmark of real GitHub issues paired with pull requests, where each task instance is embedded in a Docker-based repository snapshot with executable test suites that serve as evaluation oracles. To overcome this bottleneck, several concurrent efforts have emerged to automate large-scale environment generation. SWE-rebench (Badertdinov et al., 2025) introduces a scalable pipeline that replicates the SWE-bench construction process across a broader set of repositories, aiming to generate thousands of additional task instances with executable test environments. SWE-Universe (Chen et al., 2026b) takes a complementary approach by systematically crawling and filtering GitHub repositories to produce a diverse universe of candidate environments. SWE-Factory (Guo et al., 2026) and Scale-SWE (Zhao et al., 2026) further automate the end-to-end pipeline from repository selection to Dockerfile synthesis and test harness generation. Scale-SWE further scales this paradigm through a sandboxed multi-agent workflow. BeyondSWE (Chen et al., 2026a) expands the evaluation scope beyond single-repository bug fixing by introducing more complex real-world scenarios such as cross-repository reasoning, dependency migration, and domain-specific development tasks.SWE-World (Sun et al., 2026) proposes an orthogonal direction by replacing physical Docker execution with learned surrogate models trained on agent-environment interaction data, eliminating the resource-intensive costs of Docker environment maintenance while preserving the agent-environment feedback loop.
@@ 3.4 Dockerfile Construction
The Dockerfile agent is responsible for generating an environment for each task. During the pilot study, we identified two recurring failure modes: (1) network instability during environment construction, where generic base images require downloading Python and dependencies at build time, leading to frequent timeouts; and (2) redundant rebuilds, where unchanged base layers are reconstructed from scratch on every iteration. These inefficiencies become particularly costly at scale; therefore, we equip the Dockerfile agent with the following strategies.
@@
The Dockerfile agent will receive the repository exploration agent’s findings (e.g., special dependencies from README.md) as additional input, allowing the agent to make more informed initial decisions, and it will operate iteratively to construct the Dockerfile. If the final test execution fails, the Dockerfile agent will also receive the feedback from the test analysis agent and refine its output in subsequent attempts.
@@ 3.9 Environment Statistics
Scale-SWE (Zhao et al., 2026) 5.2k 100k 100k Real
Scale-SWE (open-sourced) 1.2k 20.2k 20.2k Real
@@ 4.1 Experimental Setup
Qwen 3 30B-A3B Series
Qwen3-30B-A3B-Instruct (Zhao et al., 2026) Qwen3-30B-A3B-Instruct OpenHands 22.0
Scale-SWE (Zhao et al., 2026) Qwen3-30B-A3B-Instruct OpenHands 64.0
@@ 4.2 Main Results
Table 3 presents the comparison of OpenSWE with representative on SWE-Bench Verified.
@@
OpenSWE-32B achieves a resolution rate of 62.4%, surpassing all methods on Qwen2.5 series. Compared to the strongest Qwen2.5-Coder-32B baseline SWE-Master-32B and SWE-Master-32B-RL. OpenSWE-32B improves by 4.6% while using a non-Coder base model, demonstrating that high-quality environment data can compensate for domain-specific pretraining. At the 72B scale, OpenSWE-72B reaches 66.0%, outperforming daVinci-Dev-72B by 7.5%. Both the 32B result and 72B result prove the effectiveness of OpenSWE.
@@ 4.4 Impact of Environment Source
To understand how the choice of environment affects downstream agent performance, we train identical models on environments from different sources and evaluate under the same protocol. Table 4 reports the results.